OpenStudy (anonymous):

what is the exact area of a circle with a radius of 31 cm A.15.5cm^2 B.31cm^2 C.961cm^2 D.240.25cm^2

OpenStudy (mathstudent55):

Do you know the formula for the area of a circle?

OpenStudy (mathstudent55):

Use it and keep \(\pi\) as \(\pi\), since they want an exact answer.

OpenStudy (anonymous):

how do i do the problem though

OpenStudy (texaschic101):

Area of a circle is : A = pi * r^2 r = 31 now just sub A = pi * 31^2

OpenStudy (anonymous):

subtract 31 from 2 ?

OpenStudy (kendricklamar2014):

If u do what @texaschic101 told you, the answer would be 31/2= 15.5 so A :)

OpenStudy (geekfromthefutur):

Kendrick already told u the answer lol its a because Area of a circle is : A = pi * r^2 r = 31 now just sub A = pi * 31^2

OpenStudy (anonymous):

yea i know i got it

OpenStudy (mathstudent55):

\(A = \pi r^2 = \pi \times (31~cm)^2 - \pi \times 31~cm \times 31~cm = 961\pi ~cm^2\)

OpenStudy (anonymous):

whats the answer yall came out whit three differ. ones what one is it???

OpenStudy (mathstudent55):

The answer is \(961 \pi ~cm^2\). None of the choices show pi, but it looks like C. is the correct answer because C. should read 961pi cm^2.